# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.2.5 Results
The first dataset is the CCMP wind speed dataset (available at http://data.remss.com), which contains 3,942 geographical grid points from the east and southeast sea areas of China. The three papers using it (Refs. [art_31, art_40, art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons. All papers present results for the complete set of 120 nodes and a subset of nodes. Since these results are similar, only the average results for all the 120 nodes are reported in Tab. LABEL:t:risultati_ccmp for the sake of simplicity.
@@
For all the time horizons considered in the papers, the dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] proposed by Gao et al. in 2023 has the highest accuracy. The DASTGN model is composed of a graph generation module and a spatio-temporal convolution module. The dynamic adaptive graph generation module generates a static graph to model long-term stable dependencies, and a dynamic graph to model the short-term dynamics. These graphs are then merged into an optimal dynamic graph. The spatio-temporal convolution module consists of an input module, a stack of spatio-temporal layers connected with residual and skip connections, and an output module that generates wind speed forecasts. Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31], and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN]. The DAGLN model was proposed by Ren at al. in 2023, and uses an encoder-decodes architecture. The encoder includes an adaptive graph learning module that learns the optimal adjacency matrix starting from a pre-defined one obtained by using the DTW algorithm, a graph convolution with node adaptive parameter learning algorithm for personalized learning of parameter space, and a GRU for temporal modeling. Then, this information is fed into the decoder that operates a convolutional and a linear operation to produce the wind speed forecast. The MTGNN model, proposed by Wu et al. in 2020, is a well-known state-of-the-art model for multivariate time series forecasting. It incorporates a graph learning layer that learns a graph adjacency matrix from scratch, followed by a stack of graph convolution and temporal convolution modules to capture spatial and temporal dependencies, and an output module. To avoid the problem of vanishing gradients, residual and skip connections are added between spatial and temporal convolutional modules.
The second dataset shared by two papers (Refs. [art_1, art_5]) is called China National Urban Air Quality dataset (available at https://github.com/Friger/GAGNN). It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples. Both these papers integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/). The average accuracy of the models used in the two papers is compared in Tab. LABEL:t:risultati_chinese_aqi.
@@
As from Tab. LABEL:t:risultati_chinese_aqi, the hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except the 1 hour horizon, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN]. The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, whereas the MegaCRN model is more focused on local features. The INNGNN model, proposed by Ding et al. in 2023, consists of a temporal and a spatial dependency module. The temporal module employs self-attention and interpretable neural networks (INN) to extract local and global features of time series data, effectively identifying periodicity and trend of the AQI time series. After that, the spatial module has an encoder-decoder architecture where GNN modules extract spatial dependencies across geographic locations, represented by an adjacency matrix based on spatial distance. The MegaCRN model, introduced by Jiang et al. in 2023 for traffic forecasting, includes a meta-graph learner module and an encoder-decoder spatial module built using a series of convolutional GNNs. The source code for the MegaCRN model can be accessed at https://github.com/deepkashiwa20/MegaCRN.
@@ 5.3 Finance
“Finance” is another emerging theme of study, though not yet widely explored, with just 16 out of 281 selected papers. The difficulty of predicting financial data is a well-established fact. This is due to the inherently complex nature of markets, the influence of geopolitical events on them, and the unpredictable behavior of humans which is often irrational and difficult to predict. This complexity makes financial time series highly volatile, and accurate forecasts and data modeling are essential to develop effective trading strategies. Spatio-temporal GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
@@ 5.3.1 Overview
The “Finance” thematic group is relatively narrow in terms of sub-themes. This is in part due to the limited number of published papers. The majority of studies are concentrated on stock price forecasting [art_59, art_60, art_62, art_63, art_65, art_377, art_382], with the aim of forecasting the future trend of the price of a company. Sometimes the stock prediction problem is seen as a classification task. This is because, due to the stochastic nature and volatility of financial markets, predicting just the direction of price movements (“up”, “neutral”, “down”) can be sometimes more feasible than accurately forecasting price values. In this case, assuming a daily granularity, a rise (“up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a “down”. Otherwise, it is classified as “neutral”. All the selected GNN papers addressing stock prices indeed approach the price forecasting problem from a classification perspective. Another distinctive aspect in this group of papers is that researchers usually evaluate, among others, how their predictions affect the stability of investment returns over a specified period. This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns [art_59, art_60, art_65, art_66, art_377]. These simulations require modeling rather than forecasting.
An examination of the selected papers indicates an interest in the application of spatio-temporal GNNs in Finance which starts from 2022. Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals. Most of the corresponding authors are affiliated with academic institutions in China.
@@ 5.3.2 Datasets
China A-Shares [art_63], [art_65], [art_66], [art_377], [art_378] https://finance.sina.com.cn/stock
Criptocurriences [art_381] https://coinmarketcap.com/
CSI 100 index [art_62] https://tushare.pro/
CSI 300 index [art_59], [art_62], [art_382] https://global.csmar.com/
Gold and silver market [art_374] https://www.kitco.com/
HAL agriculture [art_379] www.hal.gov.tr
S&P 100 index [art_62] https://finance.yahoo.com/
S&P 500 index [art_59] https://finance.yahoo.com/
Russell 1K index [art_62] https://finance.yahoo.com/
Russell 3K index [art_62] https://finance.yahoo.com/
US crude oil market [art_374] http://www.eia.gov/
All the listed datasets represent collections of stock indices. The most common datasets include the China Securities Index (CSI), the 100 and 500 Standard and Poor’s (S&P) indices, and the 100 and 300 China A-Shares indices. Many studies also include exogenous data, such as other market data, events, and news. News can be processed by some language model like BERT (as in Refs. [art_59, art_66, art_378]). Most of the data have a granularity of 1 day, and the forecasting horizon is also generally 1 day. Two of the selected papers (Refs. [art_62, art_377]) analyses longer time horizons for stock forecasting, and others do the same for agricultural price (Ref. [art_379]) or cryptocurrencies’ price (Ref. [art_381]). As for pre-processing techniques, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
@@ 5.3.3 Types of models
In terms of the taxonomy of the proposed GNN models, 6 out of 16 papers utilize attentional GNNs, 9 papers employ convolutional GNNs, and one paper takes into account both convolutional and attentional GNNs (Ref. [art_374]). As for the graph structure, some papers use models which learn the graph structure themselves (e.g., Refs. [art_62, art_64, art_378, art_379, art_382]), whereas some others define it a priori, based on the relationships between companies and stocks (e.g., Refs. [art_63, art_67, art_68, art_381]).
Half of the papers deal with a stock movement classification problem, so the most commonly used loss function is cross entropy. The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch, and a few utilizing TensorFlow. Only Refs. [art_63] and [art_379] provide a link to the source code of the proposed models, which are, respectively, the knowledge graph with graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System, and the spectral temporal graph neural network (StemGNN), available at https://github.com/cevher/Time-Series-Analysis.
@@ 5.3.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_68], [art_375]
Mean reversion (MR) [art_63], [art_65], [art_377]
Non-GNN machine learning methods Convolutional neural network (CNN) [art_59], [art_64], [art_374], [art_379]
Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65], [art_377]
Feed-forward neural network (FNN) [art_59], [art_68], [art_375], [art_378]
Gated recurrent unit (GRU) [art_60], [art_68], [art_382]
Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68], [art_374], [art_375], [art_377], [art_379], [art_381]
Support vector machine (SVM) [art_64], [art_68], [art_375], [art_378]
GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68], [art_382]
Graph convolutional network with long short term memory network (GCN+LSTM) [art_65], [art_377]
Financial Graph Attention Network (FinGAT) [FinGAT] [art_60], [art_62]
Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65], [art_377]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_62], [art_382]
Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65], [art_377]
It is evident that there is no consensus regarding the most appropriate benchmark model to use, especially of GNN type. The limited number of GNN-based benchmarks can be attributed to the fact that GNN models have not been extensively explored in the financial domain. The most common benchmark is the LSTM network, which is considered an effective approach for processing long term information in time series. The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al. in 2019 [HATS]. The HATS model was proposed for predicting stock movements using relational data. It aggregates information from different types of relationships among data and incorporates this information into each stock representation. From an architectural point of view, the model starts with employing LSTM and GRU networks as feature extraction modules. Next, a hierarchical GNN-based architecture is designed to capture the significance of neighboring nodes and relation type across different layers. Finally, the node representations are fed into a task-specific module, which can be implemented as a simple linear transformation layer. The source code for the HATS model is available at https://github.com/dmis-lab/hats.
@@ 5.3.5 Results
The accuracy of the proposed classification models is primarily evaluated in terms of accuracy, precision and F1 score. The only two papers with results which can be compared are Refs. [art_63] and [art_65], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock). These two papers compare the results obtained by using 5, 10 and 20 trading days as records. For sake of simplicity, a comparison of the accuracy of the models in the two papers with 20 days only is provided in Tab. LABEL:t:risultati_a-share.
@@
The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [art_63] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [art_65], depending on the metrics considered. The KG-GCN model, proposed by Wang et al. in 2023, first constructs a knowledge graph to represent the semantic relationships and quantify correlations between stocks. Then a community detection module based on GCN node embeddings and K-means clustering is used to group similar stocks together. Finally, the historical prices of similar stocks are fed into LSTM and GRU networks to predict stock price trends. The MONEY model, introduced by Sun et al. in 2023, is composed of a convolutional GNN and an hypergraph convolution network to capture complex relations among similar stocks.
@@ 5.4 Health
The fourth thematic group of papers is “Health”, which includes 15 out of 281 publications. These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease. All the papers working with functional magnetic resonance imaging (fMRI) data were excluded from the analysis, as the dataset was generated as image sequences rather than time series. This key difference in data structure makes them less relevant to the spatio-temporal focus of this review.
@@ 5.4.1 Overview
Applications of GNNs to the “Health” thematic group have been studied starting from 2021. The reason of this expansion may be attributed to two key factors. The first factor is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals, which can be used for the timely diagnosis of chronic diseases. The second factor is the spread of the COVID-19 pandemic, which motivated researchers to investigate and to forecast the spread of the virus. Epidemic forecasting is indeed the most common topic among the selected papers. The prediction of epidemics presents two significant challenges. First, epidemics’ seasonality can vary greatly over time, and they can manifest with different intensities and durations. Second, occasional pandemics can disrupt established patterns even for years. These complexities make the forecasting task more difficult, and non-GNN machine learning, and GNN models themselves, have been studied to address this issue. Moreover, the spread of large-scale human diseases is closely related to both the mobility patterns of populations and the demographic characteristics of residents in different regions [art_413]. Consequently, the application of GNN models to epidemic dynamics appears to be a suitable approach for analyzing and addressing such data. A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to the state of the health of individuals. The three letters of the acronym stand for susceptible (S), infectious (I), and recovered (R). Namely, the SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune. It uses a system of differential equations to describe how the disease spreads spatially and changes in time. In general, some papers use it as benchmark (e.g., Ref. [art_49, art_413, art_415]).
Other papers in this section focus on different types of time series, such as electroencephalogram (EEG) data for epilepsy detection (Ref. [art_14]) or emotion recognition (Refs. [art_411, art_412]). While emotion recognition may not seem directly related to the “Health” topic, it is included here because it relies on EEG data. The use of EEG in emotion recognition helps to identify more accurately true emotional states, as there physiological signals cannot be consciously controlled in the same way as facial expressions, tone of voice, or body posture [art_411].
The selected papers were each published in different journals, with one paper per journal, primarily in the field of computer science. The majority of the corresponding authors are affiliated with academic institutions based in China.

### v4

@@ 5.2.5 Results
The first dataset is the CCMP wind speed dataset (available at http://data.remss.com), which contains 3,942 geographical grid points from the east and southeast sea areas of China. The three papers using it (Refs. [326, 102, 97]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons. All papers present results for the complete set of 120 nodes and a subset of nodes. Since these results are similar, only the average results for all the 120 nodes are reported in Tab. LABEL:t:risultati_ccmp for the sake of simplicity.
@@
For all the time horizons considered in the papers, the dynamic adaptive spatio-temporal graph neural network (DASTGN) model [97] proposed by Gao et al. in 2023 has the highest accuracy. The DASTGN model is composed of a graph generation module and a spatio-temporal convolution module. The dynamic adaptive graph generation module generates a static graph to model long-term stable dependencies, and a dynamic graph to model the short-term dynamics. These graphs are then merged into an optimal dynamic graph. The spatio-temporal convolution module consists of an input module, a stack of spatio-temporal layers connected with residual and skip connections, and an output module that generates wind speed forecasts. Other two well performing models are the data-based adaptive graph learning network (DAGLN) [326], and the multivariate time series forecasting with graph neural network (MTGNN) [424]. The DAGLN model was proposed by Ren at al. in 2023, and uses an encoder-decodes architecture. The encoder includes an adaptive graph learning module that learns the optimal adjacency matrix starting from a pre-defined one obtained by using the DTW algorithm, a graph convolution with node adaptive parameter learning algorithm for personalized learning of parameter space, and a GRU for temporal modeling. Then, this information is fed into the decoder that operates a convolutional and a linear operation to produce the wind speed forecast. The MTGNN model, proposed by Wu et al. in 2020, is a well-known state-of-the-art model for multivariate time series forecasting, here included in the “Generic” group of papers. It incorporates a graph learning layer that learns a graph adjacency matrix from scratch, followed by a stack of graph convolution and temporal convolution modules to capture spatial and temporal dependencies, and an output module. To avoid the problem of vanishing gradients, residual and skip connections are added between spatial and temporal convolutional modules.
The second dataset shared by two papers (Refs. [41, 72]) is called China National Urban Air Quality dataset (available at https://github.com/Friger/GAGNN). It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples. Both these papers integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/). The average accuracy of the models used in the two papers is compared in Tab. LABEL:t:risultati_chinese_aqi.
@@
As from Tab. LABEL:t:risultati_chinese_aqi, the hybrid interpretable neural network and graph neural network model (INNGNN) [72] is the most accurate model for all time horizons except the 1 hour horizon, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [166]. The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, whereas the MegaCRN model is more focused on local features. The INNGNN model, proposed by Ding et al. in 2023, consists of a temporal and a spatial dependency module. The temporal module employs self-attention and interpretable neural networks (INN) to extract local and global features of time series data, effectively identifying periodicity and trend of the AQI time series. After that, the spatial module has an encoder-decoder architecture where GNN modules extract spatial dependencies across geographic locations, represented by an adjacency matrix based on spatial distance. The MegaCRN model, introduced by Jiang et al. in 2023 for traffic forecasting, includes a meta-graph learner module and an encoder-decoder spatial module built using a series of convolutional GNNs. The source code for the MegaCRN model can be accessed at https://github.com/deepkashiwa20/MegaCRN.
@@ 5.3 Finance
“Finance” is another emerging theme of study, though not yet widely explored, with just 14 out of 366 selected papers. The difficulty of predicting financial data is a well-established fact. This is due to the inherently complex nature of markets, the influence of geopolitical events on them, and the unpredictable behavior of humans which is often irrational and difficult to predict. This complexity makes financial time series highly volatile, and accurate forecasts and data modeling are essential to develop effective trading strategies. Spatio-temporal GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
@@ 5.3.1 Overview
The “Finance” thematic group is relatively narrow in terms of sub-themes. This is in part due to the limited number of published papers. Half of the studies are concentrated on stock price forecasting [151, 307, 371, 391, 362, 37, 429], with the aim of forecasting the future trend of the price of a company. Sometimes the stock prediction problem is seen as a classification task. This is because, due to the stochastic nature and volatility of financial markets, predicting just the direction of price movements (“up”, “neutral”, “down”) can be sometimes more feasible than accurately forecasting price values. In this case, assuming a daily granularity, a rise (“up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a “down”. Otherwise, it is classified as “neutral”. All the selected GNN papers addressing stock prices indeed approach the price forecasting problem from a classification perspective. Another distinctive aspect in this group of papers is that researchers usually evaluate, among others, how their predictions affect the stability of investment returns over a specified period. This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns [151, 307, 362, 53, 429]. These simulations require modeling rather than forecasting.
An examination of the selected papers indicates an interest in the application of spatio-temporal GNNs in Finance which starts from 2022. Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals. The only two conference papers of this group were presented at CIKM.
@@ 5.3.2 Datasets
China A-Shares [391], [362], [53], [205] https://finance.sina.com.cn/stock
CSI 100 index [371] https://tushare.pro/
CSI 300 index [151], [371], [37], [429] https://global.csmar.com/
Nordpool dataset [343] https://www.nordpoolgroup.com/
S&P 100 index [371] https://finance.yahoo.com/
S&P 500 index [151], [429] https://finance.yahoo.com/
Russell 1K index [371] https://finance.yahoo.com/
Russell 3K index [371] https://finance.yahoo.com/
All the listed datasets represent collections of stock indices. The most common datasets include the China Securities Index (CSI), the 100 and 500 Standard and Poor’s (S&P) indices, and the 100 and 300 China A-Shares indices. Many studies also include exogenous data, such as other market data, events, and news. News can be processed by some language model like BERT (as in Refs. [151, 53, 205]). Most of the data have a granularity of 1 day, and the forecasting horizon is also generally 1 day. One of the selected papers (Ref. [371]) analyses longer time horizons for stock forecasting. As for pre-processing techniques, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
@@ 5.3.3 Types of models
In terms of the taxonomy of the proposed GNN models, 7 out of 14 papers utilize attentional GNNs, 6 papers employ convolutional GNNs, and another one (Ref. [343]) does not allow us to specify the classification of the model within the proposed taxonomy. As for the graph structure, some papers use models which learn the graph structure themselves (e.g., Refs. [371, 225, 205, 37]), whereas some others define it a priori, based on the relationships between companies and stocks (e.g., Refs. [391, 158, 444]). One paper (Ref. [448]) examines how graph structures change across different observation scales, based on the idea that variable correlations observed over short timescales may differ significantly from those seen over longer periods.
Half of the papers deal with a stock movement classification problem, so the most commonly used loss function is cross entropy. The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch, and a few utilizing TensorFlow. Only Ref. [391] provides a link to the source code of the proposed models, which is the knowledge graph with graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System.
@@ 5.3.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [444], [297]
Mean reversion (MR) [391], [362]
Non-GNN machine learning methods Convolutional neural network (CNN) [151], [225]
Dual-stage attention-based recurrent neural network (DARNN) [391], [362]
Feed-forward neural network (FNN) [151], [444], [297], [205]
Gated recurrent unit (GRU) [307], [444], [37], [429]
Long-short term memory network (LSTM) [151], [371], [391], [225], [362], [158], [444], [297], [429]
Support vector machine (SVM) [225], [444], [297], [205]
Transformer [375] [37], [429]
GNN methods Graph convolutional network (GCN) [151], [307], [371], [225], [444], [37]
Financial graph attention network (FinGAT) [142] [307], [371]
Hierarchical graph attention network (HATS) [188] [371], [391], [362], [429]
Multivariate time series forecasting with graph neural network (MTGNN) [424] [371], [37]
Spatiotemporal hypergraph convolution network (STHGCN) [331] [391], [362]
Temporal graph convolution (TGC) [88] [391], [429]
It is evident that there is no consensus regarding the most appropriate benchmark model to use, especially of GNN type. The limited number of GNN-based benchmarks can be attributed to the fact that GNN models have not been extensively explored in the financial domain. The most common benchmark is the LSTM network, which is considered an effective approach for processing long term information in time series. The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al. in 2019 [188]. The HATS model was proposed for predicting stock movements using relational data. It aggregates information from different types of relationships among data and incorporates this information into each stock representation. From an architectural point of view, the model starts with employing LSTM and GRU networks as feature extraction modules. Next, a hierarchical GNN-based architecture is designed to capture the significance of neighboring nodes and relation type across different layers. Finally, the node representations are fed into a task-specific module, which can be implemented as a simple linear transformation layer. The source code for the HATS model is available at https://github.com/dmis-lab/hats.
@@ 5.3.5 Results
The accuracy of the proposed classification models is primarily evaluated in terms of accuracy, precision and F1 score. The only two papers with results which can be compared are Refs. [391] and [362], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock). These two papers compare the results obtained by using 5, 10 and 20 trading days as records. For sake of simplicity, a comparison of the accuracy of the models in the two papers with 20 days only is provided in Tab. LABEL:t:risultati_a-share.
@@
The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [391] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [362], depending on the metrics considered. The KG-GCN model, proposed by Wang et al. in 2023, first constructs a knowledge graph to represent the semantic relationships and quantify correlations between stocks. Then a community detection module based on GCN node embeddings and K-means clustering is used to group similar stocks together. Finally, the historical prices of similar stocks are fed into LSTM and GRU networks to predict stock price trends. The MONEY model, introduced by Sun et al. in 2023, is composed of a convolutional GNN and an hypergraph convolution network to capture complex relations among similar stocks.
@@ 5.4 Health
The fourth thematic group of papers is “Health”, which includes 23 out of 366 publications. These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease. All the papers working with functional magnetic resonance imaging (fMRI) data were excluded from the analysis, as the dataset was generated as image sequences rather than time series. This key difference in data structure makes them less relevant to the spatio-temporal focus of this review.
@@ 5.4.1 Overview
Applications of GNNs to the “Health” thematic group have been studied starting from 2021. The reason of this expansion may be attributed to two key factors. The first factor is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals, which can be used for the timely diagnosis of chronic diseases. The second factor is the spread of the COVID-19 pandemic, which motivated researchers to investigate and to forecast the spread of the virus. Epidemic forecasting is indeed the most common topic among the selected papers. The prediction of epidemics presents two significant challenges. First, epidemics’ seasonality can vary greatly over time, and they can manifest with different intensities and durations. Second, occasional pandemics can disrupt established patterns even for years. These complexities make the forecasting task more difficult, and non-GNN machine learning, and GNN models themselves, have been studied to address this issue. Moreover, the spread of large-scale human diseases is closely related to both the mobility patterns of populations and the demographic characteristics of residents in different regions [445]. Consequently, the application of GNN models to epidemic dynamics appears to be a suitable approach for analyzing and addressing such data. A common reference model for epidemic prediction is the SIR model [180], an acronym referring to the state of the health of individuals. The three letters of the acronym stand for susceptible (S), infectious (I), and recovered (R). Namely, the SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune. It uses a system of differential equations to describe how the disease spreads spatially and changes in time. The main advantage of this model is its inherent interpretability; it produces coefficients that are easy to interpret. However, its strong and simplistic assumptions also restrict its forecasting capability [514]. In general, some papers use it as benchmark (e.g., Ref. [462, 445, 285]).
One study (Ref. [334]) also analyzes how air quality and satellite observations of temperature and humidity relate to COVID-19 mortality and severity.
Other papers in this section focus on different types of time series, such as electroencephalogram (EEG) data for epilepsy detection (Refs. [383, 38, 23]) or emotion recognition (Refs. [527, 119, 384]). While emotion recognition may not seem directly related to the “Health” topic, it is included here because it relies on EEG data. The use of EEG in emotion recognition helps to identify more accurately true emotional states, as there physiological signals cannot be consciously controlled in the same way as facial expressions, tone of voice, or body posture [527].
The selected papers were each published in different journals, with one paper per journal, primarily in the field of computer science. In terms of conference venues, KDD and CIKM are the two most popular, with three papers each.